OpenStudy (anonymous):

Simplify each of the following expressions by applying one of the theorems. (a) X′Y′Z + (X′Y′Z)′ (b) (AB′+ CD)(B′E + CD) (c) ACF + AC′F (d) A(C + D′B) + A′ (e) (A′B + C + D)(A′B + D) (f) (A + BC)+(DE + F)(A + BC)′

OpenStudy (ajprincess):

Does this come under boolean algebra?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

please help me to answer this

OpenStudy (ajprincess):

What is A+A'=?

OpenStudy (anonymous):

1

OpenStudy (ajprincess):

In the first question (X'Y'Z)' is the complement of (X'Y'Z) do u see any relation between that theorem and ur first question?

OpenStudy (anonymous):

how will i know if there is a relation between a theorem?

OpenStudy (ajprincess):

A+A'=1 right? If so what is assume (X'Y'Z) as A (X'Y'Z)+(X'Y'Z)'=?

OpenStudy (anonymous):

1

OpenStudy (ajprincess):

yes:)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

but i need to show a solution in my paper how i get that answer.

OpenStudy (ajprincess):

u can state the theorem u used to get the answer. A+A'=1 It's complement law isnt it?

OpenStudy (anonymous):

yes it is

OpenStudy (ajprincess):

so u can write by complement law (X'Y'Z)+(X'Y'Z)'=1

OpenStudy (anonymous):

thanks. can you help me with the other questions?

OpenStudy (ajprincess):

ya. For the second u need remove the brackets. for that u have to use distributive law. (AB′+ CD)(B′E + CD)=AB'(B'E+CD)+CD(BE'+CD)

OpenStudy (ajprincess):

AB'(B'E+CD)+CD(BE'+CD)=?

OpenStudy (anonymous):

X(Y+Z)=XY+XZ <<-- this ?

OpenStudy (ajprincess):

yes. using that expand AB'(B'E+CD)+CD(BE'+CD

OpenStudy (anonymous):

i dont know how. im soory

OpenStudy (ajprincess):

AB'(B'E+CD)+CD(BE'+CD=AB'*B'E+AB'*CD+CD*BE'+CD*CD

OpenStudy (ajprincess):

nw see if u can use this law in that A*A=A

OpenStudy (anonymous):

sorry if im to slow to reply. coz im trying to solve what law u give :)

OpenStudy (ajprincess):

that's k:)

OpenStudy (anonymous):

are you busy? coz im having a trouble solving it. :(

OpenStudy (ajprincess):

nope i am free. take ur time and tell me where u r stuck>

OpenStudy (anonymous):

i dont know w/c i will use in

OpenStudy (anonymous):

the law that u give to me.

OpenStudy (anonymous):

how i will i know what will be the given in A*A?

OpenStudy (ajprincess):

AB'(B'E+CD)+CD(BE'+CD=AB'*B'E+AB'*CD+CD*BE'+CD*CD Did u get this?

OpenStudy (ajprincess):

in CD*CD u have CD instead of A that u had in A*A. following me?

OpenStudy (anonymous):

how did u get BE'? IS THAT B'E?

OpenStudy (ajprincess):

oops sorry. yes it is B'E. am really sorry

OpenStudy (anonymous):

no worries. :) lets proceed

OpenStudy (ajprincess):

ha k:) in CD*CD u have CD instead of A that u had in A*A. so if A*A=A then CD*CD=CD. following me?

OpenStudy (anonymous):

yah

OpenStudy (ajprincess):

AB'(B'E+CD)+CD(B'E+CD=AB'*B'E+AB'*CD+CD*B'E+CD*CD =AB'B'E+AB'CD+CDB'E+CD is this k?

OpenStudy (anonymous):

ohhhh.. i get it :)

OpenStudy (ajprincess):

consider CDB'E+CD can u factor out anything?

OpenStudy (anonymous):

cd?

OpenStudy (ajprincess):

ya. factor it out and simplify

OpenStudy (anonymous):

CDB'E+CD in this given i will factor out and simplify?

OpenStudy (ajprincess):

ya. CDB'E+CD factor out CD(B'E+1) A+1=1 so CD(B'E+1)=CD*1=CD.getting it?

OpenStudy (anonymous):

yah

OpenStudy (ajprincess):

AB'B'E+AB'CD+CDB'E+CD=AB'B'E+AB'CD+CD =AB'E+CD

OpenStudy (anonymous):

how will i know if i will use the laws?

OpenStudy (ajprincess):

AB'CD+CD check again. if u can factor out anything. yes. u can factor CD Then when u factor out CD u will get CD(AB'+1) u knw A+1=1 u have AB' instead of A. So AB'+1=1 Then CD(AB'+1)=CD AB'B'E as u can see B'B' is similar to AA=A so B'B'=B so AB'B'E=AB'E. so finally AB'B'E+AB'CD+CD=AB'E+CD

OpenStudy (anonymous):

ok. gets

OpenStudy (ajprincess):

hw abt trying this nw? ACF + AC′F

OpenStudy (anonymous):

try only . but not sure.

OpenStudy (ajprincess):

tell me ur answer

OpenStudy (anonymous):

can u start first then i will follow?

OpenStudy (ajprincess):

try factor out the common thing in it

OpenStudy (anonymous):

af(?????)

OpenStudy (ajprincess):

yup

OpenStudy (anonymous):

uhmm. im having a trouble with the c' and c. i dont know what will i write

OpenStudy (ajprincess):

ACF+AC'F=AF(C+C')

OpenStudy (ajprincess):

(C+C') does this look similar to any laws u knw?

OpenStudy (anonymous):

1?

OpenStudy (ajprincess):

yup:) so what is AF(C+C')

OpenStudy (anonymous):

af

OpenStudy (ajprincess):

yup good:)

OpenStudy (anonymous):

hehe.

OpenStudy (anonymous):

(A′B + C + D)(A′B + D)=A'B(A'B+D)+C(A'B+D)+D(A'B+D) is this right??

OpenStudy (ajprincess):

yup:)

OpenStudy (ajprincess):

expand it and simplify:)

OpenStudy (anonymous):

=a'b*a'b+a'b*d+c*a'b+c*d+d*a'b+d*d =a'ba'b+a'bd+c'ab+cd+da'b+d =a'ba'b+a'bd+ca'b+cd+d am i right?

OpenStudy (ajprincess):

yup:) u can simplify it further

OpenStudy (anonymous):

uhm. thats my problem now. hehe.

OpenStudy (ajprincess):

a'ba'b=?

OpenStudy (anonymous):

a'b

OpenStudy (ajprincess):

yup:)

OpenStudy (ajprincess):

cd+d=?

OpenStudy (anonymous):

d?

OpenStudy (ajprincess):

yup:)

OpenStudy (ajprincess):

a'ba'b+a'bd+ca'b+cd+d=?

OpenStudy (anonymous):

a'b+d ???hehe

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

but what happen to the a'bd+ca'b?

OpenStudy (ajprincess):

a'ba'b+a'bd+ca'b+cd+d=a'b+a'bd+ca'b+d =a'b(1+d)+ca'b+d =a'b+ca'b+d =a'b(1+c)+d =a'b+d

OpenStudy (anonymous):

can you still help me with the last 2 questions?

OpenStudy (ajprincess):

let me try them out and tell u

OpenStudy (anonymous):

ok. thanks

OpenStudy (rational):

(e) (A′B + C + D)(A′B + D) ( A′B + D + C)(A′B + D) ------- -------- say A'B + D = X (X + C)X XX + CX X + CX X(1+C) X A'B + D

OpenStudy (rational):

(f) (A + BC)+(DE + F)(A + BC)' ------- -------- say A+BC = X, DE + F = Y X + X'Y (X+X')(X+Y) 1(X+Y) X + Y A + BC + DE + F

OpenStudy (ajprincess):

A(C+D'B)+A'=AC+AD'B+A' =AC+AD'B+A'.1 (C+1=1) =AC+AD'B+A'(C+1) =AC+AD'B+A'C+A' =C(A+A')+AD'B+A'.1 =C+AD'B+A'(DB'+1) (D'B+1=1) =C+AD'B+A'D'B+A' =C+D'B(A+A')+A' =C+D'B+A'

OpenStudy (rational):

another way for D : (d) A(C + D′B) + A′ say X = A', C+D'B = Y X'Y + X (X'+X)(X+Y) X + Y A' + C + D'B

OpenStudy (anonymous):

hi rational im a bit having a trouble with the last part of the solution u had. but it is good coz it is short way to solve.

OpenStudy (rational):

below distributive law is a powerful property : \(x+yz = (x+y)(x+z)\)

OpenStudy (anonymous):

hi ajprincess can u give me another way to solve letter f?

OpenStudy (rational):

e and f uses that property, so if you know how to use that property there should not be any problems

OpenStudy (ajprincess):

I totally forgot about that law. thanx a lot @rational. i remembered it nw cos of u.:) well alternative way to solve f would be the one i used for d part. bt I guess it will be too lengthy.

OpenStudy (anonymous):

thanks @ajprincess i learned a lot. hope that u can help me again. @rational thanks for solving alternative way to solve :)

OpenStudy (ajprincess):

U r welcome:) glad to knw that u learnt a lot :)

OpenStudy (anonymous):

hi @ajprincess may i ask a question. in this part >>> =AC+AD'B+A'.1 (how you get A'.1?)

OpenStudy (ajprincess):

when we multiply a number or variable by 1 the answer is the same as number or variable. i.e A*1=A 3*1=3 I used the converse part